OpenStudy (anonymous):

how do i find the slope and y intercept of a line using its equation 7y+12x-2=0

OpenStudy (anonymous):

best way is to solve for y

OpenStudy (anonymous):

you get \[7y=-12x+2\] \[y=-\frac{12}{7}x+\frac{2}{7}\]

OpenStudy (anonymous):

which looks like \[y=mx+b\] so you can read off your answer

OpenStudy (anonymous):

You can turn this into slope intercept form 7y+12x-2 =0 7y+12x= 2 7y = -12x+2 7y/7 = -12x/7+2/7 y = -12/7 + 2/7